The dimensions of a rectangle are such that it's length is 9in more than it's width. If the length were doubled and if the width were decreased by 4in, the area would be increased by 168in squared. What are the length and width of the rectangle?
Answer by jorel1380(2518)

l=w+9
2l(w-4)-168=lw
2(w+9)(w-4)-168=(w+9)(w)
2(w2+5w-36)-168=w2+9w
2w2+10w-72-168=w2+9w
w2+w-240=0
(w+16)(w-15)=0
w=15 or -16
Throwing out the negative answer, we get the width to be 15 in., and the length to be 24 in..